# Understanding uncertainty in Bayesian cluster analysis

### v1

@@ Abstract
The Bayesian approach to clustering is often appreciated for its ability to provide uncertainty in the partition structure. However, summarizing the posterior distribution over the clustering structure can be challenging, due the discrete, unordered nature and massive dimension of the space. While recent advancements provide a single clustering estimate to represent the posterior, this ignores uncertainty and may even be unrepresentative in instances where the posterior is multimodal. To enhance our understanding of uncertainty, we propose a WASserstein Approximation for Bayesian clusterIng (WASABI), which summarizes the posterior samples with not one, but multiple clustering estimates, each corresponding to a different part of the space of partitions that receives substantial posterior mass. Specifically, we find such clustering estimates by approximating the posterior distribution in a Wasserstein distance sense, equipped with a suitable metric on the partition space. An interesting byproduct is that a locally optimal solution to this problem can be found using a k-medoids-like algorithm on the partition space to divide the posterior samples into different groups, each represented by one of the clustering estimates. Using both synthetic and real datasets, we show that our proposal helps to improve the understanding of uncertainty, particularly when the data clusters are not well separated or when the employed model is misspecified.
@@ 1 Introduction
Clustering is one of the canonical forms of unsupervised learning and has numerous applications in various contexts, from finance and economics (Saunders, 1980), to natural language processing and computer science (Blei et al., 2003), biomedicine (Oyelade et al., 2016; Chauvel et al., 2019) and more. It aims to discover groups of similar data points and characterize the patterns within. While algorithmic schemes, such as k-means (Hartigan and Wong, 1979; Jain, 2010), are widely used, only a single clustering solution is provided, without any measure of uncertainty in the solution. To overcome this, the Bayesian approach is a natural choice. Through Bayes rule, the prior over the space of clusterings is combined with the model (or loss) to assess homogeneity within group, producing posterior beliefs and uncertainty for all possible clustering structures and patterns within each cluster.
@@
However, understanding this posterior over the clustering structure can be daunting, due to the huge dimension and discrete, unordered nature of the space of partitions. Indeed, the number of clusters is rarely known in practice, and many Bayesian methods account for this uncertainty, further increasing the dimension of the space. For example, in a Bayesian nonparametric setting, allowing anywhere between 1 and $n$ clusters, the total number of ways to partition the data is a Bell number, which is $B_{20}=51,724,158,235,372$ for only 20 data points. Even if the number of clusters is restricted in a parametric setting, the number of possible partitions is still massive, e.g. restricting to at most 5 clusters, there are still $795,019,337,135$ possible clusterings of $20$ data points.
@@
Clearly, this makes any exact posterior computations infeasible, and to overcome this, Markov chain Monte Carlo (MCMC) is the most common tool. MCMC provides tens of thousands of clustering solutions, representing (asymptotically exact) draws from the posterior. Due to high dimension of the space, almost all of these clustering solutions are unique and describing the behavior and uncertainty across them is far from straightforward. Alternative inference procedures can also be used, such as importance sampling (Zhang and Williamson, 2019), sequential Monte Carlo (MacEachern et al., 1999; Härkönen et al., 2022), or posterior bootstrap (Fong et al., 2019), which similarly provide a very large number of different clustering solutions. An increasingly used alternative to posterior simulation is approximate optimization-based techniques, such as variational inference. However, variational inference is well known to suffer from mode collapse, with the approximate posterior concentrating around a single clustering solution. Thus, it tends to provide unsatisfying and unrealistic measures of certainty in the clustering structure. A simple approach to overcome this combines multiple variational approximations through ensembles or stacking (Yao et al., 2022), which similarly produces a large number of different clustering solutions. Ensembles are also used to combat poor mixing of MCMC in high-dimensions through consensus clustering (Coleman et al., 2022).
@@
When faced with this challenge of describing the numerous clustering solutions produced by these inference algorithms, a first task is to find a single representative clustering. From a decision-theoretic viewpoint, the optimal Bayes estimator of the clustering is found by minimizing the posterior expected loss. This requires the choice of an appropriate loss on the space of partitions, and the 0-1 loss is a simple choice, resulting in the maximum a posteriori (MAP) partition estimator. However, the 0-1 loss between two partitions is zero if and only if the two partitions coincide, and it equally penalizes any two different partitions, irrespective of how similar or different they are. This property makes the 0-1 loss quite unappealing. Thus, more general losses are required and have been used, such as Binder’s loss (Binder, 1978; Dahl, 2006; Lau and Green, 2007), the adjusted Rand Index (ARI, Fritsch and Ickstadt, 2009), the variation of information (VI, Meilă, 2007; Wade and Ghahramani, 2018), the normalized VI (Vinh et al., 2010; Rastelli and Friel, 2018), the generalized Binder’s loss and VI (Dahl et al., 2022), among others (Quintana and Iglesias, 2003; Buch et al., 2024; Nguyen and Mueller, 2024). However, these methods all focus on finding a single optimal clustering estimate (a critical aspect that is also highlighted in the discussion of (Rastelli and Friel, 2018; Dahl et al., 2022)).
@@
Beyond providing a single clustering solution, a main motivation for the Bayesian approach is the additional uncertainty in the clustering structure. Thus, tools are also required to visualize and describe this uncertainty, which again is difficult due to the sheer number of partitions, combined with the variable number of clusters and label-switching. One important tool is the posterior similarity matrix, whose elements represent the posterior probability that a pair of data points are clustered together (which is also referred to as the co-clustering or consensus matrix). Additionally, Wade and Ghahramani (2018) developed credible balls to characterize uncertainty around the clustering estimate, but accurate estimation of the bounds of the credible ball based on the MCMC draws can be challenging in certain cases due to massive dimension of the partition space. An alternative greedy approach to compute the credible bounds, that allows searching beyond the MCMC draws, is presented in Buch et al. (2024). Another approach (Lavigne and Liverani, 2024) approximates and visualizes uncertainty in the allocation of each data point, but only provides a limited view of uncertainty by conditioning on the allocation of all other points. Lastly, Do et al. (2024) represent uncertainty in the underlying mixing measure for Gaussian mixtures through a dendogram.
(a) Posterior on partitions
(b) Posterior similarity matrix
(c) Histogram of the data
(d) WASABI summarizes with multiple, weighted clustering estimates
@@
Figure 1: Slightly bimodal example. The MAP partition of a DPM for data generated from a slightly bimodal mixture (Figure 1(c)) contains only a single cluster, yet the posterior similarity matrix (Figure 1(b)) suggests uncertainty in additional clusters. To further examine the posterior on the space of partitions, Figure 1(a) shows the log posterior relative to the MAP against the VI distance to the MAP, where each point represents a partition colored by its number of clusters, with cluster sizes reported for some. Multimodality in the posterior on partitions is evident with one mode corresponding to the MAP and another around partitions with two clusters of more equal size. WASABI summarizes with multiple, weighted partitions (Figure 1(d)), reflecting the two different modes of clustering in this example.
@@
In this work, we contribute an important novel tool to enhance our understanding of uncertainty in the posterior on partitions, that is applicable to the wide range of Bayesian clustering models for diverse data types. As an illustrative toy example, consider data generated from a slightly bimodal mixture of Gaussians (Figure 1(c)). Interestingly, when focusing on the marginal posterior of the number of clusters, a number of papers Miller and Harrison (2013, 2014); Ohn and Lin (2020); Ascolani et al. (2022); Alamichel et al. (2024) have demonstrated inconsistency and overestimation of the number of clusters in Bayesian nonparametric mixture models, such as the popular Dirichlet process mixture (DPM), if the hyperparameters are not appropriately tuned. Despite this, the optimal clustering estimate can still recover the true clustering structure under an appropriate choice of loss (see (Rajkowski, 2019) for a theoretical analysis of the MAP and Wade (2023) for an empirical analysis comparing the MAP, VI and Binder’s loss). However, when the clusters are not well separated, the different estimators often disagree; indeed the MAP and VI clustering estimate only contain a single cluster [^1] While previous work found agreement between the clustering solutions which minimize the expected VI and its more computationally efficient lower bound, in this slightly bimodal example, they disagree, with the latter reporting two clusters., while the ARI and Binder’s contain two clusters in the slightly bimodal example (Figure 1(c)). The posterior similarity matrix (Figure 1(b)) also suggests possible additional clusters, but without a clear separation between groups. Further examination of the posterior on partitions (Figure 1(a)) shows multimodality in the posterior. Starting from the MAP partition with a single cluster, the posterior mass decreases slightly as a single data point is split off, and continues to decrease as we move further from the MAP, creating additional small clusters. But, it starts to increase again around more distant partitions with two clusters of roughly equal size.
@@
To capture this multimodality and describe the main clustering patterns present across the posterior, we propose to summarize the posterior with not one, but multiple optimal partitions, found by minimizing the Wasserstein distance equipped with an appropriate loss. Our WASserstein Approximation for Bayesian clusterIng (WASABI) also provides weights attached to the optimal partitions (referred to as particles) and reflects the different modes of clustering in the slightly bimodal example (Figure 1(d)). We focus on the VI loss, due to its benefits shown in (e.g Wade and Ghahramani, 2018; Rastelli and Friel, 2018; Dahl et al., 2022), but other losses can also be considered. This builds on the idea of Balocchi et al. (2023), who proposed an ensemble optimization method that approximates the posterior distribution by targeting several partition estimates through minimization of the Kullback-Leibler divergence. Instead, we use the Wasserstein distance, motivated by its role in hierarchical modeling and clustering. Indeed, it has been shown to be the natural distance metric for convergence theory of mixing measures (Nguyen, 2013, 2016; Ho and Nguyen, 2016; Guha et al., 2021). Furthermore, the Wasserstein distance is one of the few distances that allows measuring the discrepancy between distributions based on an arbitrary metric on the space on which they are supported. An interesting byproduct of using the Wasserstein distance is that the problem of determining the optimal discrete distribution with minimal Wasserstein distance from the empirical distribution can be directly linked to k-mediods clustering (Pollard, 1982; Graf and Luschgy, 2000; Canas and Rosasco, 2012). As a consequence, the tool we propose can be seen as employing k-mediods clustering to divide the samples into distinct groups and using the mediods as optimal partitions summarizing the multiple modes of the posterior. This will greatly enhance researchers’ toolbox to better describe and communicate findings and showcase the benefits of the uncertainty provided by the Bayesian approach to clustering. The code implementing the WASABI algorithm is available as an R package at https://github.com/cecilia-balocchi/WASABI.
@@ 2 Review
In the context of clustering, the observed data consists of measurements $\bm{y}=(y_{1},\ldots,y_{n})$ drawn from a heterogeneous population consisting of an unknown number homogeneous sub-populations. The observed $y_{i}\in\mathcal{Y}$ may be continuous, discrete, mixed, or more complex in nature (e.g. functional data). Each data point is associated with a discrete latent variable $c_{i}$ (also called the allocation variable) indicating the group membership of the data point, i.e. $c_{i}=j$ if $y_{i}$ belongs to the $j$th group, and $c_{i}=c_{i^{\prime}}$ if $y_{i}$ and $y_{i^{\prime}}$ belong to the same sub-population. We are interested in obtaining estimates and understanding the uncertainty of clustering structure characterized by the latent $\rho_{n}=(c_{1},\ldots,c_{n})$. To achieve this, the Bayesian approach constructs a posterior distribution over the space of partitions,
@@
where $\pi(\rho_{n})$ represents the prior over the space of partitions and $p(\bm{y}\mid\rho_{n})$ is the likelihood in a model-based approach or defined based on a loss function in a loss-based approach (Rigon et al., 2020). For reviews on Bayesian clustering, we refer the reader to Wade (2023); Grazian (2023).
It is worth emphasizing that clustering is often referred to as an ill-posed problem, as it aims to discover unknown patterns or structures in the data. The notion of a cluster depends on the application at hand and can often be challenging to characterize formally. A unique clustering solution often does not exist (Hennig, 2015). Thus, one must carefully consider the model or loss employed and importantly, also characterize uncertainty in the clustering solution. To achieve the latter, Bayesian cluster analysis provides a formal framework through both the posterior distribution over the entire space of clusterings and by creating an ensemble of clustering solutions sampled from the posterior. Moreover, this also helps to mitigate sensitivity to local optimum which adversely impact all clustering algorithms due to the sheer size of the space.
@@ 2.1 Random Partition Models
A key ingredient in the Bayesian approach is the prior $\pi(\rho_{n})$ over the clustering structure. This is referred to as random partition model, as it involves randomly assigning observations to clusters, or equivalently, randomly partitioning the indices of the data points $\{1,\ldots,n\}$ into $K_{n}$ non-empty and mutually exhaustive sets $C_{j}$ for $j=1,\ldots,K_{n}$, where $K_{n}$ represents the number of clusters. Thus, the partition $\rho_{n}$ can also be represented as $\rho_{n}=\{C_{1},\ldots,C_{K_{n}}\}$, where the sets satisfy $C_{j}\cap C_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $C_{1}\cup\ldots\cup C_{K_{n}}=\{1,\ldots,n\}$; essentially, each $C_{j}$ contains the indices of data points in the $j$th cluster. For ease of notation, the sample size is dropped from $\rho_{n}$ and $K_{n}$, when the context is clear.
In practice, the number of clusters $K$ is rarely known, and choosing this number is an important and difficult concern in cluster analysis. While information criteria or model selection tools, such as the Bayesian information criterion (BIC), are commonly used, this ignores uncertainty in the number of clusters and disregards information from the numerous models fits using different choices of $K$. Instead, our focus is approaches that account for this uncertainty, namely, through 1) mixtures of finite mixtures, which naturally incorporate a prior on the unknown $K$ Miller and Harrison (2018); Nobile and Fearnside (2007); Richardson and Green (1997); 2) sparse overfitted mixtures that specify an upper bound on the number of clusters and encourage extra components to be emptied through sparsity promoting priors (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Van Havre et al., 2015; Frühwirth-Schnatter et al., 2021); and 3) Bayesian nonparametric (BNP) mixtures (Müller, 2019) that allow the number of clusters to grow unboundedly with the data, with the Dirichlet process (DP) mixture (Lo, 1984) being the most-widely used example. For example, letting $L$ denote an upper bound on the number of clusters and $\gamma>0$, the random partition model for the overfitted mixture (corresponding to a Dirichlet prior on the weights of the mixture with $L$ components) is

### v2

# Interpreting uncertainty in Bayesian cluster analysis

@@ Abstract
The Bayesian approach to clustering is often appreciated for its ability to provide uncertainty in the partition structure. However, summarizing the posterior distribution over the clustering structure can be challenging, due the discrete, unordered nature and massive dimension of the space. While recent advancements provide a single clustering estimate to represent the posterior, this ignores uncertainty and may even be unrepresentative in instances where the posterior is multimodal. To enhance our understanding of uncertainty, we propose a WASserstein Approximation for Bayesian clusterIng (WASABI), which summarizes the posterior samples with not one, but multiple clustering estimates, each corresponding to a different part of the partition space that receives substantial posterior mass. Specifically, we find such clustering estimates by approximating the posterior distribution in a Wasserstein distance sense, equipped with a suitable metric on the partition space. An interesting byproduct is that a locally optimal solution can be found using a k-medoids-like algorithm on the partition space to divide the posterior samples into groups, each represented by one of the clustering estimates. Using synthetic and real datasets, we show that WASABI helps to improve the understanding of uncertainty, particularly when clusters are not well separated or when the employed model is misspecified.
@@ 1 Introduction
Clustering is one of the canonical forms of unsupervised learning and has numerous applications in various contexts, from finance and economics (Saunders, 1980), to natural language processing and computer science (Blei et al., 2003), biomedicine (Oyelade et al., 2016) and more. It aims to discover groups of similar data points and characterize the patterns within. While algorithmic schemes, such as k-means (Hartigan and Wong, 1979), are widely used, only a single clustering solution is provided, without any measure of uncertainty in the solution. To overcome this, the Bayesian approach is a natural choice. Through Bayes rule, the prior over the space of clusterings is combined with the model (or loss) to assess homogeneity within group, producing posterior beliefs and uncertainty for all possible clustering structures and patterns within each cluster.
@@
However, interpreting this posterior over the clustering structure can be daunting, due to the huge dimension and discrete, unordered nature of the space of partitions. Indeed, the number of clusters is rarely known in practice, and many Bayesian methods account for this uncertainty, further increasing the dimension of the space. For example, in a Bayesian nonparametric (BNP) setting, allowing anywhere between 1 and $n$ clusters, the total number of ways to partition the data is a Bell number, which is $B_{20}=51,724,158,235,372$ for only 20 data points. Even if the number of clusters is restricted in a parametric setting, the number of possible partitions is still massive, e.g. restricting to at most 5 clusters, there are still $795,019,337,135$ possible clusterings of $20$ data points.
@@
Clearly, this makes any exact posterior computations infeasible, and to overcome this, Markov chain Monte Carlo (MCMC) is the most common tool. MCMC provides tens of thousands of clustering solutions, representing (asymptotically exact) draws from the posterior. Due to high dimension of the space, almost all of these clustering solutions are unique and describing the behavior and uncertainty across them is far from straightforward. Alternative inference procedures can also be used, such as importance sampling (Zhang and Williamson, 2019), sequential Monte Carlo (MacEachern et al., 1999), or posterior bootstrap (Fong et al., 2019), which similarly provide a very large number of different clustering solutions. An increasingly used alternative to posterior simulation is approximate optimization-based techniques, such as variational inference (VI). However, VI is known to suffer from mode collapse, with the approximate posterior concentrating around a single clustering solution and providing unrealistic measures of certainty. An approach to overcome this combines multiple variational approximations through ensembles or stacking (Yao et al., 2022), which similarly produces a large number of different clustering solutions. Ensembles are also used to combat poor mixing of MCMC in high-dimensions through consensus clustering (Coleman et al., 2022).
@@
When faced with the challenge of describing the numerous clustering solutions produced by these inference algorithms, a first task is to find a single representative clustering. From a decision-theoretic view, the optimal Bayes estimator of the clustering is found by minimizing the posterior expected loss. This requires the choice of an appropriate loss on the space of partitions, and the 0-1 loss is a simple choice, resulting in the maximum a posteriori (MAP) partition estimator. However, the 0-1 loss between two partitions is zero if and only if they coincide, and it equally penalizes any two different partitions, irrespective of how similar or different they are. This property makes the 0-1 loss unappealing. Thus, more general losses are required and have been used, such as Binder’s loss (Binder, 1978; Dahl, 2006; Lau and Green, 2007), the adjusted Rand Index (ARI, Fritsch and Ickstadt, 2009), the variation of information (VI, Meilă, 2007; Wade and Ghahramani, 2018), the normalized VI (Vinh et al., 2010; Rastelli and Friel, 2018), the generalized Binder’s loss and VI (Dahl et al., 2022), among others (Quintana and Iglesias, 2003; Franzolini and Rebaudo, 2024; Buch et al., 2024; Nguyen and Müller, 2024). However, these methods focus on a single optimal clustering estimate (a critical aspect also highlighted in the discussion of Rastelli and Friel (2018); Dahl et al. (2022)).
@@
Beyond providing a single clustering solution, a main motivation for the Bayesian approach is the additional uncertainty in the clustering structure. Thus, tools are also required to visualize and describe this uncertainty, which again is difficult due to the sheer number of partitions, combined with the variable number of clusters and label-switching. One important tool is the posterior similarity matrix (PSM), whose elements represent the posterior probability that a pair of data points are clustered together (also referred to as the co-clustering or consensus matrix). Additionally, Wade and Ghahramani (2018) developed credible balls to characterize uncertainty around the clustering estimate, but accurate estimation of the bounds of the credible ball based on the MCMC draws can be challenging in certain cases due to massive dimension of the partition space. An alternative greedy approach to compute the credible bounds, that allows searching beyond the MCMC draws, is presented in Buch et al. (2024). Another approach (Lavigne and Liverani, 2024) approximates and visualizes uncertainty in the allocation of each data point, but only provides a limited view of uncertainty by conditioning on the allocation of all other points. Lastly, Do et al. (2024) represent uncertainty in the underlying mixing measure for Gaussian mixtures through a dendogram.
(a) Histogram of the data
(b) Posterior on partitions
(c) Posterior similarity matrix
(d) WASABI summarizes with multiple clustering estimates
@@
Figure 1: Slightly bimodal example. The MAP partition of a DPM for data generated from a slightly bimodal mixture (Figure 1(a)) contains only a single cluster, yet the posterior similarity matrix (Figure 1(c)) suggests uncertainty in additional clusters. To further examine the posterior on the space of partitions, Figure 1(b) shows the log posterior relative to the MAP against the VI distance to the MAP, where each point represents a partition colored by its number of clusters, with cluster sizes reported for some (the set of partitions plotted consists of the MAP and those with 2 clusters that respect the order of the observed data). Multimodality in the posterior is evident with one mode corresponding to the MAP and another around partitions with two clusters of more equal size. WASABI summarizes with multiple, weighted partitions (Figure 1(d)), reflecting the two different modes of clustering in this example.
@@
In this work, we contribute an important novel tool to enhance our understanding of uncertainty in the posterior on partitions, that is applicable to the wide range of Bayesian clustering models for diverse data types. As an illustrative toy example, consider data generated from a slightly bimodal mixture of Gaussians (Figure 1(a)). A common paradox in BNP mixture models, such as the popular Dirichlet process mixture (DPM, Lo, 1984), is that the marginal posterior on the number of clusters may suffer from inconsistency and overestimation (Miller and Harrison, 2014; Ohn and Lin, 2023; Ascolani et al., 2023; Alamichel et al., 2024), if hyperparameters are not appropriately tuned; yet, the optimal clustering estimate can still recover the true structure under appropriate choice of loss (see Rajkowski (2019) for a theoretical analysis of the MAP and Wade (2023) for an empirical analysis of the MAP, VI, and Binder). However, when the clusters are not well separated, Rajkowski (2019) shows that the MAP underestimates the number of clusters. Indeed, in the slightly bimodal example, the MAP and VI estimate only contain a single cluster, while the ARI and Binder overestimate with 4 and 59 clusters, respectively (see Section D.1 of the Appendix). The posterior similarity matrix (Figure 1(c)) suggests possible additional clusters and further examination of the posterior on partitions (Figure 1(b)) shows multimodality. Starting from the MAP partition with a single cluster, the posterior mass decreases slightly when a single data point is split off, and continues to decrease as we increase the size of the small cluster and move further from the MAP. But, it starts to increase for more distant partitions with two clusters of more equal size.
@@
To capture this multimodality and describe the main clustering patterns present across the posterior, we propose to summarize the posterior with not one, but multiple optimal partitions, found by minimizing the Wasserstein distance equipped with an appropriate loss. Our WASserstein Approximation for Bayesian clusterIng (WASABI) also provides weights attached to the optimal partitions (referred to as particles) and reflects the different modes of clustering in the slightly bimodal example (Figure 1(d)). We focus on the VI loss, due to its benefits shown in Wade and Ghahramani (2018); Rastelli and Friel (2018); Dahl et al. (2022), but other losses can be considered. This builds on the idea of Balocchi et al. (2023), who proposed an ensemble optimization method that approximates the posterior by targeting several partition estimates through minimization of the Kullback-Leibler divergence. Instead, we use the Wasserstein distance, motivated by its role in hierarchical modeling and clustering. Indeed, it has been shown to be the natural distance metric for convergence theory of mixing measures (Nguyen, 2013; Ho and Nguyen, 2016; Guha et al., 2021). Furthermore, the Wasserstein distance is one of the few distances that allows measuring the discrepancy between distributions based on an arbitrary metric on the space on which they are supported. An interesting byproduct of using the Wasserstein distance is that the problem of determining the optimal discrete distribution with minimal Wasserstein distance from the empirical distribution can be directly linked to k-medoids clustering (Pollard, 1982; Canas and Rosasco, 2012). As a consequence, the tool we propose can be seen as employing k-medoids clustering to divide the samples into distinct groups and using the medoids as optimal partitions summarizing the multiple modes of the posterior. This will greatly enhance researchers’ toolbox to better describe and communicate findings and showcase the benefits of the uncertainty provided by the Bayesian approach to clustering. The code implementing the WASABI algorithm is available as an R package at https://github.com/cecilia-balocchi/WASABI.
@@ 2 Review
In the context of clustering, the observed data consists of measurements $\bm{y}=(y_{1},\ldots,y_{n})$ drawn from a heterogeneous population consisting of an unknown number homogeneous sub-populations. The observed $y_{i}\in\mathcal{Y}$ may be continuous, discrete, mixed, or more complex in nature (e.g. functional data). Each data point is associated with a discrete latent variable $c_{i}$ (also called the allocation variable) indicating the group membership of the data point, i.e. $c_{i}=j$ if $y_{i}$ belongs to the $j$th group, and $c_{i}=c_{i^{\prime}}$ if $y_{i}$ and $y_{i^{\prime}}$ belong to the same sub-population. We are interested in obtaining estimates and understanding the uncertainty of clustering structure characterized by the latent $\rho_{n}=(c_{1},\ldots,c_{n})$. To achieve this, the Bayesian approach constructs the posterior distribution
@@
where $\pi(\rho_{n})$ represents the prior over the space of partitions and $p(\bm{y}\mid\rho_{n})$ is the likelihood in a model-based approach or defined based on a loss function in a generalized Bayes approach (Rigon et al., 2023). For reviews of Bayesian clustering, we refer the reader to Wade (2023); Grazian (2023).
It is worth emphasizing that clustering is often referred to as an ill-posed problem, as it aims to discover unknown patterns or structures in the data. The notion of a cluster depends on the application at hand and can be challenging to characterize formally. There is often no unique clustering solution (Hennig, 2015). Thus, one must carefully consider the model or loss employed and importantly, also characterize uncertainty in the clustering solution. To achieve the latter, Bayesian cluster analysis provides a formal framework through the posterior distribution over the entire space of clusterings. Moreover, this helps to mitigate sensitivity to local optima which adversely impacts all clustering algorithms due to the sheer size of the space.
@@ 2.1 Random Partition Models
A key ingredient in the Bayesian approach is the prior $\pi(\rho_{n})$ on the partition space. This is referred to as random partition model, as it involves randomly assigning observations to clusters, or equivalently, randomly partitioning the indices of the data points $\{1,\ldots,n\}$ into $K_{n}$ non-empty and mutually exhaustive sets $C_{j}$ for $j=1,\ldots,K_{n}$, where $K_{n}$ represents the number of clusters. Thus, the partition $\rho_{n}$ can also be represented as $\rho_{n}=\{C_{1},\ldots,C_{K_{n}}\}$, where the sets satisfy $C_{j}\cap C_{j^{\prime}}=\emptyset$ for $j\neq j^{\prime}$ and $C_{1}\cup\ldots\cup C_{K_{n}}=\{1,\ldots,n\}$; essentially, each $C_{j}$ contains the indices of data points in the $j$th cluster. For ease of notation, the sample size is dropped from $\rho_{n}$ and $K_{n}$, when the context is clear.
In practice, the number of clusters $K$ is rarely known, and choosing this number is an important concern in cluster analysis. While information criteria or model selection tools, such as the Bayesian information criterion (BIC), are commonly used, this ignores uncertainty in the number of clusters and disregards information from the numerous models fits using different choices of $K$. Instead, our focus is approaches that account for this uncertainty, namely, through 1) mixtures of finite mixtures, which naturally incorporate a prior on the unknown $K$ (Miller and Harrison, 2018; Nobile and Fearnside, 2007; Richardson and Green, 1997); 2) sparse overfitted mixtures that specify an upper bound on the number of clusters and empty extra components through sparsity promoting priors (Rousseau and Mengersen, 2011; Malsiner-Walli et al., 2016; Frühwirth-Schnatter et al., 2021); and 3) Bayesian nonparametric (BNP) mixtures (Müller, 2019) that allow the number of clusters to grow unboundedly with the data. For example, letting $L$ denote an upper bound on $K$ and $\gamma>0$, the random partition model for the overfitted mixture (with a Dirichlet prior on the mixture weights) is